# Fluctuation response of a superconductor with temporally correlated noise

### v1

@@ Abstract
Physical implementation.– Up to this point, we have treated Eq. (1) primarily as a toy model for a superconductor subject to colored noise. Its microscopic derivation from the BCS model inherently assumes equilibrium internal interactions of quasiparticles with phonons or impurities, which result in overdamped dissipations near $T_{c}$. In order to capture nonequilibrium fluctuation behavior, one can introduce an additional coupling to an external bath with a nonthermal occupation function $n_{b}(\omega)$. Within the Keldysh formalism, it is possible to systematically integrate out the bath and electronic degrees of freedom, and then introduce order parameter fluctuations [23, 32, 3]. In the regime $|T-T_{c}|\ll T_{c}$ this yields exactly the same coarse-grained Gaussian MSR action (3), where the noise correlator $D(\omega)$ is determined by the bath distribution $n_{b}(\omega)$. If the characteristic energy scale of the bath $\omega_{c}\sim 1/\tau$ is sufficiently low, such that $\tau\sim\tau_{\text{GL}}$, then the system can experience colored noise. The realization of the NESS with broken detailed balance in this case depends sensitively on the physical bath model. One may suggest that bosonic (and potentially fermionic) bath engineering [33, 34] allows to design specific forms of $n_{b}(\omega)$ using, for instance, external irradiation with spectral filtering [35] or out-of-equilibrium current or voltage biasing [22].
Conclusion.– We have shown that finite correlation time of Langevin noise in the stochastic TDGL model affects the electric and heat transport in fluctuating superconductors by unevenly driving order-parameter modes and suppressing high-energy contributions. This selective driving can enhance certain components of the thermoelectric response despite truncating the noise spectrum. For exponentially correlated noise, the nature and strength of this effect depends solely on the system dimensionality. This suggests a potential pathway for controlling superconducting transport properties via coupling to a nonthermal bath. Future work should explore the role of spatial correlations and the development of suitable microscopic models to connect these finding more directly with experimental platforms.
Acknowledgments.– V. P. thanks A. Varlamov, F. Piazza, A. Buzdin and M. Pini for insightful discussions and valuable comments on this work. This research is supported by the Staatsministerium für Wissenschaft und Kunst through the Hightech Agenda Bayern Plus and is part of the Munich Quantum Valley.

### v2

@@ Abstract
Physical implementation.– Up to this point, we have treated Eq. (1) primarily as a toy model for a superconductor subject to colored noise. Its microscopic derivation from the BCS model assumes equilibrium internal interactions of quasiparticles with phonons or impurities, which result in overdamped dissipations near $T_{c}$. In order to capture nonequilibrium fluctuation behavior, one can introduce an additional coupling to an external bath with a nonthermal occupation function $n_{b}(\omega)$. Within the Keldysh formalism, it is possible to systematically integrate out the bath and electronic degrees of freedom, and then introduce order parameter fluctuations [26, 36, 5]. In the regime $|T-T_{c}|\ll T_{c}$ this procedure yields exactly the same coarse-grained Gaussian MSR action (3), where the noise correlator $D(\omega)$ is determined by the bath distribution $n_{b}(\omega)$. If the characteristic energy scale of the bath $\omega_{c}\sim 1/\tau$ is sufficiently low, such that $\tau\gtrsim\tau_{\text{GL}}$, then the system can experience colored noise. The realization of the NESS with broken detailed balance in this case depends sensitively on the physical bath model and its coupling to the system. One may suggest that bosonic (and potentially fermionic) bath engineering [37, 38] allows to design specific forms of $n_{b}(\omega)$ using, for instance, external irradiation with spectral filtering [39] or out-of-equilibrium current or voltage biasing [23].
Conclusion.– We have shown that finite correlation time of Langevin noise in the stochastic TDGL model affects the charge and heat transport in fluctuating superconductors by unevenly driving order-parameter modes and suppressing high-energy contributions. This selective driving can enhance certain components of thermoelectric response despite the truncation the noise spectrum. For exponentially correlated noise, the nature and strength of this effect depends solely on the system dimensionality. These findings indicate that coupling to a nonthermal bath can provide a practical means of tuning superconducting transport. Future work should explore the role of spatial correlations and the development of suitable microscopic models to connect these finding more directly with experimental platforms.
Acknowledgements.
V. P. thanks A. Varlamov and M. Pini for insightful discussions, and F. Piazza and A. Buzdin for valuable comments. This research was supported by the Staatsministerium für Wissenschaft und Kunst through the Hightech Agenda Bayern Plus, is part of the Munich Quantum Valley, and received additional support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the Walter Benjamin Programme (Project No. 566401345).